OpenStudy (anonymous):

a projectile is thrown upward so that its distance above the ground after t seconds is h(t)=-15t^2+390t. after how many seconds does it reach its maximum height?

OpenStudy (anonymous):

A lot

OpenStudy (anonymous):

get the first derivative and equal it to zero -30t+390=0 t= 13 sec

OpenStudy (anonymous):



OpenStudy (anonymous):

@Kellbell456 did you get it

OpenStudy (anonymous):

good luck with that program...

OpenStudy (anonymous):

@satellite73 what do you mean by program??

OpenStudy (anonymous):

First you need to determine what the maximum height is.

OpenStudy (anonymous):

i bet $7 this is not a calculus class first coordinate of the vertex is \(-\frac{b}{2a}\) is what is asked for here

OpenStudy (anonymous):

\[h(t)=-15t^2+390t\] \[-\frac{b}{2a}=-\frac{390}{2\times (-15)}\] etc

OpenStudy (anonymous):

equation is wrong in any case but no matter

OpenStudy (anonymous):

13 looks good to me

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

Thank you satellite73. :)